oxidation number oxidation number for N in K3Fe(NO3)6 please explain how to find it

oxidation number



oxidation number for N in K3Fe(NO3)6
please explain how to find it
Answer #1

K3[Fe(NO3)6]  

N is present in nitrate ion  

nitrate ion = NO3-

assume N = X

O = -2 oxidation state = 3 O atoms so 3 (-2) = -6

we have to equilize -1 as -1 charge present on ion.

now

X + 2(-3) = -1

X - 6 = -1

X = -1 + 6

X = +5

oxidation state on N = +5